OpenStudy (anonymous):

solve equation (r-3) (r+6) = 0 ... someone teach me how to do these so, i don't have to keep asking questions

OpenStudy (turingtest):

the zero factor property stares that if we have a situation\[ab=0\] then either \(a=0,b=0\) ot both so we split this up and set each factor=0 on your case we can say\[a=r-3=0\]and\[b=r+6=0\]

OpenStudy (turingtest):

states*

OpenStudy (turingtest):

or both* sorry for the typos

OpenStudy (turingtest):

so if we had\[(x+9)(x-13)=0\]we would do the same thing; solve each separately\[x+9=0\]\[x-13=0\]

OpenStudy (turingtest):

this works no matter how many factors we have if there were four we could do the same thing\[x(x+2)(x-5)(x+7)=0\]means we would have to solve\[x=0\](which is already solved)\[x+2=0\]\[x-5=0\]\[x+7=0\]

OpenStudy (anonymous):

Thank you so much this helps alot

OpenStudy (turingtest):

welcome!